A 1.67kg rock is dropped from the top of a cliff that is 32.7m above the surface of a lake. What is the speed of the rock just before it hits the surface of the water
1
Expert's answer
2020-04-28T09:46:21-0400

The law of conservation of energy says

"E_i=E_f"

Hence


"mgh=\frac{mv^2}{2}"

"v=\sqrt{2gh}=\sqrt{2\times 9.8\times 32.7}=25.3\:\rm m/s"